OpenStudy (anonymous):

Find the length of the altitude of an equilateral triangle if one side has a length of 6 cm

OpenStudy (anonymous):

\[\sin 60 \times 6\]

OpenStudy (anonymous):

does it come out too 5.196152423 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok thank you

OpenStudy (anonymous):

wait my answers dont have that at all A 3√5 cm B 6√2 cm C 3√3 cm

OpenStudy (anonymous):

C)\[3\sqrt{3}\]

OpenStudy (anonymous):

Thank you